# Implications of recent Higgs data for the light Higgs bosons in the GNMSSM

### v2

@@ 3.2 HiggsTools constraints
The branching ratios for the exotic decays $h\to a_{1}a_{1}$ and $h\to h_{1}h_{1}$ depend on their respective triple Higgs couplings $C_{ha_{1}a_{1}}$ and $C_{hh_{1}h_{1}}$, as defined in Eqs. (2.14) and (2.15). These couplings are primarily governed by the parameters $\lambda$, $\kappa$, and the SM-like fraction $V_{h}^{\rm SM}$ of the $h$ boson. In the $h_{2}$ scenario, $C_{hh_{1}h_{1}}$ is additionally modulated by the singlet fraction $V_{h_{s}}^{\rm S}$ of the singlet-dominated state $h_{s}$. Figure 4 displays the correlation between $\lambda$ and $\kappa$ under HiggsTools constraints, indicating $\lambda\leq 0.1$ for most viable parameter points, which are consistent with experimental preferences for minimal doublet-singlet mixing. Further, Figs. 5 and 6 reveal the Higgs boson compositions. Both $h_{1}$ and $h_{2}$ scenarios require the observed $h$ boson to be predominantly SM-like ($V_{h}^{\rm SM}\geq 0.93$) with suppressed singlet admixture ($V_{h}^{\rm S}\leq 0.32$). For the $h_{2}$ scenario, the lighter $h_{s}$ state must exhibit high singlet purity ($V_{h_{s}}^{\rm S}\geq 0.94$) and negligible SM-like components ($V_{h_{s}}^{\rm SM}\leq 0.32$).
@@ 4 Conclusion
This study systematically investigates exotic decays of the 125 GeV Higgs boson into pairs of the lightest CP-odd or CP-even Higgs bosons within the GNMSSM. Our analysis incorporates the latest constraints from ATLAS direct searches and utilizes the integrated framework of HiggsTools, which combines updated exclusion data from HiggsBounds and statistical analyses from HiggsSignals.
We first scanned the GNMSSM parameter space under fundamental constraints imposed by HiggsSignals-2.6.2 and HiggsBounds-5.10.2. The HiggsTools framework was then applied to exhaustively constrain this parameter space. A comprehensive comparison reveals that HiggsTools imposes significantly stronger constraints than previous individual versions of HiggsSignals and HiggsBounds. Furthermore, its exclusion power substantially surpasses current experimental limits derived from ATLAS direct Higgs searches, which are solely based on individual decay channel analyses.
To elucidate the impact of recent Higgs data on GNMSSM scenarios featuring a light Higgs boson, we separately analyzed the constraining effects of the subpackages HiggsSignals and HiggsBounds within HiggsTools, leading to the following key conclusions:
@@
  • In both the $h_{1}$ and $h_{2}$ scenarios, constraints from HiggsBounds prove dominant in restricting the theoretical parameter space. Constraints from HiggsSignals and direct ATLAS searches targeting individual decay channels are comparatively weaker. Nearly all parameter points excluded by HiggsSignals or individual LHC channel analyses are also excluded by HiggsBounds. This stems from HiggsBounds providing direct constraints for non-SM Higgs searches, whereas HiggsSignals imposes indirect constraints.
@@
  • Analyses within the $h_{2}$ scenario reveal distinct phenomenological features. Notably, HiggsSignals can exclude parameter points exhibiting suppressed exotic decay branching ratios, specifically $Br(h\to a_{1}a_{1}\to\tau\tau bb)\leq 2.5\%$. This exclusion arises due to the kinematic accessibility of the decay $h_{2}\to h_{1}h_{1}$, which features significantly enhanced branching fractions. These augmented decays modify the properties of the SM-like Higgs boson, thereby increasing the sensitivity of HiggsSignals’ indirect constraints in otherwise kinematically suppressed regions.
@@
  • Under the combined constraints of HiggsTools, both the $h_{1}$ and $h_{2}$ scenarios require the observed $h$ boson to be predominantly SM-like ($V_{h}^{\rm SM}\geq 0.93$) with minimal singlet admixture ($V_{h}^{\rm S}\leq 0.32$). In the $h_{2}$ scenario, the lighter $h_{s}$ state must exhibit high singlet purity ($V_{h_{s}}^{\rm S}\geq 0.94$) and negligible SM-like components ($V_{h_{s}}^{\rm SM}\leq 0.32$).
Crucially, despite the rigorous constraints imposed by current Higgs data on light Higgs scenarios, phenomenologically viable regions of parameter space persist. These regions offer promising prospects for future exploration of light Higgs phenomenology.

### v3

@@ 3.2 HiggsTools constraints
The branching ratios for the exotic decays of $h\to a_{1}a_{1}$ and $h\to h_{1}h_{1}$ depend on their respective triple Higgs couplings $C_{ha_{1}a_{1}}$ and $C_{hh_{1}h_{1}}$, as defined in Eqs. (2.18) and (2.19). These couplings are primarily governed by parameters $\lambda$ and $\kappa$ and the SM-like fraction $V_{h}^{\rm SM}$ of the $h$ boson. In the $h_{2}$ scenario, $C_{hh_{1}h_{1}}$ is also modulated by the singlet fraction $V_{h_{s}}^{\rm S}$ of the singlet-dominated state $h_{s}$. Figure 4 displays the correlation between $\lambda$ and $\kappa$ under HiggsTools constraints, which indicates that $\lambda\leq 0.1$ for most viable parameter points, which is consistent with experimental preferences for minimal doublet-singlet mixing. Figures 5 and 6 reveal the Higgs boson compositions. Both $h_{1}$ and $h_{2}$ scenarios require that the observed $h$ boson is predominantly SM-like ($V_{h}^{\rm SM}\geq 0.93$) with suppressed singlet admixture ($V_{h}^{\rm S}\leq 0.32$). For the $h_{2}$ scenario, the lighter $h_{s}$ state must exhibit high singlet purity ($V_{h_{s}}^{\rm S}\geq 0.94$) and negligible SM-like components ($V_{h_{s}}^{\rm SM}\leq 0.32$). Figure 6 further illustrates that $V_{h}^{S}\simeq-V_{h_{s}}^{\rm SM}$, confirming the theoretical derivation presented in Eq.(2.14).
Point P1 Point P2
Parameter Value Parameter Value
$\lambda$, $\kappa$, 0.015,0.3690.015,-0.369 $\lambda$, $\kappa$ 0.017,0.4550.017,-0.455
$\tan\beta$, $\delta$ $41.5,0.02$ $\tan\beta$, $\delta$ $37.6,0.08$
$v_{s}$, $\mu_{\rm tot}$, $m_{N}$ $212.8,468.9,23.6$ $v_{s}$, $\mu_{\rm tot}$, $m_{N}$ $811.7,342.0,147.7$
$m_{B}$, $m_{C}$ $208.8,144.6$ $m_{B}$$m_{C}$ $181.6,23.3$
$A_{t}$, $A_{\kappa}$ 2536.3,685.32536.3,-685.3 $A_{t}$$A_{\kappa}$ 2592.3,1152.82592.3,-1152.8
Particle Mass Spectrum Particle Mass Spectrum
$\tilde{\chi}_{1}^{0}$$\tilde{\chi}_{2}^{0}$, $\tilde{\chi}_{3}^{0}$ $23.5,478.2,483.5$ $\tilde{\chi}_{1}^{0}$$\tilde{\chi}_{2}^{0}$, $\tilde{\chi}_{3}^{0}$ $146.5,349.2,354.2$
$\tilde{\chi}_{4}^{0}$, $\tilde{\chi}_{5}^{0}$ $1008.3,2020.8$ $\tilde{\chi}_{4}^{0}$, $\tilde{\chi}_{5}^{0}$ $1008.1,2020.8$
$\tilde{\chi}_{1}^{\pm}$, $\tilde{\chi}_{2}^{\pm}$ $481.4,2021.1$ $\tilde{\chi}_{1}^{\pm}$, $\tilde{\chi}_{2}^{\pm}$ $352.2,2021.0$
$_{s}, hh$, $H$ $154.8,125.1,2414.2$ $_{s}, hh$, $H$ $188.9,125.3,2273.6$
$A_{s}$, $A_{H}$ $48.0,2414.1$ $A_{s}$, $A_{H}$ $16.6,2273.6$
Primary Decays Branching Ratios [%] Primary Decays Branching Ratios [%]
$h\to b\bar{b}$ $47.5$ $h\to b\bar{b}$ $49.4$
$h\to W^{+}W^{-}$ $22.4$ $h\to W^{+}W^{-}$ $24.0$
$h\to A_{s}A_{s}$ $16.1$ $h\to A_{s}A_{s}$ $11.9$
Rotation Matrix Element Value Rotation Matrix Element Value
$V_{h_{s}}^{S}$$V_{h_{s}}^{SM}$$V_{h}^{S}$$V_{h}^{SM}$ 0.99996, 0.0005,0.0003,0.9997-0.99996,\ 0.0005,\ -0.0003,\ -0.9997 $V_{h_{s}}^{S}$$V_{h_{s}}^{SM}$$V_{h}^{S}$$V_{h}^{SM}$ 0.99996,0.005, 0.005,0.9996-0.99996,\ -0.005,\ 0.005,\ -0.9996
$N_{11}$$N_{12}$$N_{13}$$N_{14}$$N_{15}$ 0.0003,0.0002,0.0002,0.0055,0.999980.0003,-0.0002,0.0002,-0.0055,0.99998 $N_{11}$$N_{12}$$N_{13}$$N_{14}$$N_{15}$ 0.00059, 0.0004,0.004, 0.00997,0.9999-0.00059,\ 0.0004,\ -0.004,\ 0.00997,\ -0.9999
$N_{21}$$N_{22}$$N_{23}$$N_{24}$$N_{25}$ -0.0599, 0.036, -0.71, 0.70, 0.004 $N_{21}$$N_{22}$$N_{23}$$N_{24}$$N_{25}$ 0.048, -0.033, 0.71, -0.70, -0.00996
$N_{31}$$N_{32}$$N_{33}$$N_{34}$$N_{35}$ 0.02, -0.02, -0.71, -0.71, -0.0037 $N_{31}$$N_{32}$$N_{33}$$N_{34}$$N_{35}$ -0.022, 0.023, 0.70, 0.71, 0.004
$N_{41}$$N_{42}$$N_{43}$$N_{44}$$N_{45}$ 0.998, 0.0045, -0.028, 0.056, 0 $N_{41}$$N_{42}$$N_{43}$$N_{44}$$N_{45}$ -0.998, -0.004, 0.02, -0.05, 0
$N_{51}$$N_{52}$$N_{53}$$N_{54}$$N_{55}$ 0.0019, -0.999, -0.01, 0.04, 0 $N_{51}$$N_{52}$$N_{53}$$N_{54}$$N_{55}$ 0.002, -0.999, -0.0078, 0.04, 0
Primary Annihilation Fraction [%] Primary Annihilation Fraction [%]
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to b\bar{b}$ $79.97$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}A_{s}$ $81.2$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to\tau^{+}\tau^{-}$ $19.74$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to A_{s}A_{s}$ $18.79$
DM Observable Value DM Observable Value
$\Omega h^{2}$ $0.000086$ $\Omega h^{2}$ $0.001593$
$\sigma^{\rm SI}_{p,n}/(10^{-49}\text{cm}^{2})$ $6.8,7.9$ $\sigma^{\rm SI}_{p,n}/(10^{-48}\text{cm}^{2})$ $4.1,4.8$
$\sigma^{\rm SD}_{p,n}/(10^{-50}\text{cm}^{2})$ $2.5,1.9$ $\sigma^{\rm SD}_{p,n}/(10^{-48}\text{cm}^{2})$ $3.6,2.8$
Table 2: Benchmark points P1 and P2 compatible with all experimental constraints. Both points correspond to Singlino-dominated DM in the $h_{1}$ scenario. All mass parameters are given in GeV.
Point P3 Point P4
Parameter Value Parameter Value
$\lambda$, $\kappa$, $0.16,0.117$ $\lambda$, $\kappa$ 0.026,0.0160.026,-0.016
$\tan\beta$, $\delta$ 20.4,0.0220.4,-0.02 $\tan\beta$, $\delta$ $23.7,0.108$
$v_{s}$, $\mu_{\rm tot}$, $m_{N}$ $739.4,269.5,776.5$ $v_{s}$, $\mu_{\rm tot}$, $m_{N}$ 872.6,405.6,655.4872.6,405.6,-655.4
$m_{B}$, $m_{C}$ $282.3,240.0$ $m_{B}$$m_{C}$ $51.2,73.5$
$A_{t}$, $A_{\kappa}$ 2694.0,1683.32694.0,-1683.3 $A_{t}$$A_{\kappa}$ $2720.2,1859.9$
Particle Mass Spectrum Particle Mass Spectrum
$\tilde{\chi}_{1}^{0}$$\tilde{\chi}_{2}^{0}$, $\tilde{\chi}_{3}^{0}$ $273.8,279.6,777.0$ $\tilde{\chi}_{1}^{0}$$\tilde{\chi}_{2}^{0}$, $\tilde{\chi}_{3}^{0}$ $413.0,418.1,655.4$
$\tilde{\chi}_{4}^{0}$, $\tilde{\chi}_{5}^{0}$ $1008.1,2020.8$ $\tilde{\chi}_{4}^{0}$, $\tilde{\chi}_{5}^{0}$ $1008.2,2020.7$
$\tilde{\chi}_{1}^{\pm}$, $\tilde{\chi}_{2}^{\pm}$ $277.3,2021.0$ $\tilde{\chi}_{1}^{\pm}$, $\tilde{\chi}_{2}^{\pm}$ $416.1,2021.0$
$_{s}, hh$, $H$ $169.1,125.3,2104.7$ $_{s}, hh$, $H$ $21.77,125.4,2177.9$
$A_{s}$, $A_{H}$ $61.2,2103.9$ $A_{s}$, $A_{H}$ $56.8,2177.9$
Primary Decays Branching Ratios [%] Primary Decays Branching Ratios [%]
$h\to b\bar{b}$ $52.1$ $h\to b\bar{b}$ $48.4$
$h\to W^{+}W^{-}$ $24.0$ $h\to W^{+}W^{-}$ $23.7$
$h\to A_{s}A_{s}/h_{s}h_{s}$ $8.8/0$ $h\to A_{s}A_{s}/h_{s}h_{s}$ $3.6/9.96$
Rotation Matrix Element Value Rotation Matrix Element Value
$V_{h_{s}}^{S}$$V_{h_{s}}^{SM}$$V_{h}^{S}$$V_{h}^{SM}$ 0.999, 0.03,0.03,0.998-0.999,\ 0.03,\ -0.03,\ -0.998 $V_{h_{s}}^{S}$$V_{h_{s}}^{SM}$$V_{h}^{S}$$V_{h}^{SM}$ 0.9997,0.021, 0.021,0.99890.9997,\ -0.021,\ 0.021,\-0.9989
$N_{11}$$N_{12}$$N_{13}$$N_{14}$$N_{15}$ 0.044,0.033,0.71,0.70,0.0370.044,-0.033,0.71,-0.70,0.037 $N_{11}$$N_{12}$$N_{13}$$N_{14}$$N_{15}$ 0.054,0.035, 0.71,0.70,0.00290.054,\ -0.035,\ 0.71,\ -0.70,\ -0.0029
$N_{21}$$N_{22}$$N_{23}$$N_{24}$$N_{25}$ 0.023, 0.023, 0.70, 0.71, 0.019-0.023,\ 0.023,\ 0.70,\ 0.71,\ 0.019 $N_{21}$$N_{22}$$N_{23}$$N_{24}$$N_{25}$ 0.021,0.022,$0.71,$ 0.0140.021,\ -0.022,\ -0.71,\ -0.71,\ 0.014
$N_{31}$$N_{32}$$N_{33}$$N_{34}$$N_{35}$ 0.0001, 0.0009,0.04, 0.01, 0.999-0.0001,\ 0.0009,\ -0.04,\ 0.01,\ 0.999 $N_{31}$$N_{32}$$N_{33}$$N_{34}$$N_{35}$ 0.0001, 0.0002, 0.012, 0.0078, 0.999899-0.0001,\ 0.0002,\ 0.012,\ 0.0078,\ 0.999899
$N_{41}$$N_{42}$$N_{43}$$N_{44}$$N_{45}$ 0.9987,0.004, 0.015,0.047, 0.001-0.9987,\ -0.004,\ 0.015,\ -0.047,\ 0.001 $N_{41}$$N_{42}$$N_{43}$$N_{44}$$N_{45}$ 0.998,0.004, 0.02,0.05, 0-0.998,\ -0.004,\ 0.02,\ -0.05,\ 0
$N_{51}$$N_{52}$$N_{53}$$N_{54}$$N_{55}$ 0.002, 0.999, 0.007,0.039,0.0001-0.002,\ 0.999,\ 0.007,\ -0.039,\ -0.0001 $N_{51}$$N_{52}$$N_{53}$$N_{54}$$N_{55}$ 0.002,0.999,0.0098, 0.04, 00.002,\ -0.999,\ -0.0098,\ 0.04,\ 0
Primary Annihilation Fraction [%] Primary Annihilation Fraction [%]
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to d\bar{u}/\bar{d}u$ $10.58$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to d\bar{u}/\bar{d}u$ $10.05$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to s\bar{c}/\bar{s}c$ $10.58$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to s\bar{c}/\bar{s}c$ $10.05$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to b\bar{t}/t\bar{b}$ $7.8$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to b\bar{t}/t\bar{b}$ $6.66$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to W^{+}W^{-}$ $6.7$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to W^{+}W^{-}$ $4.84$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to ZZ$ $5.21$ $\tilde{\chi}_{2}^{0}\tilde{\chi}_{1}^{\pm}\to d\bar{u}/\bar{d}u$ $4.41$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to e\bar{\nu_{e}}/\bar{\nu_{e}}e$ $3.53$ $\tilde{\chi}_{2}^{0}\tilde{\chi}_{1}^{\pm}\to s\bar{c}/\bar{s}c$ $4.41$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to\mu\bar{\nu_{\mu}}/\bar{\nu_{\mu}}\mu$ $3.53$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to ZZ$ $3.81$
$\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to\tau\bar{\nu_{\tau}}/\bar{\nu_{\tau}}\mu$ $3.50$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to e\bar{\nu_{e}}/\bar{\nu_{e}}e$ $3.35$
$\tilde{\chi}_{2}^{0}\tilde{\chi}_{1}^{\pm}\to d\bar{u}/\bar{d}u$ $3.17$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to\mu\bar{\nu_{\mu}}/\bar{\nu_{\mu}}\mu$ $3.35$
$\tilde{\chi}_{2}^{0}\tilde{\chi}_{1}^{\pm}\to s\bar{c}/\bar{s}c$ $3.17$ $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{\pm}\to\tau\bar{\nu_{\tau}}/\bar{\nu_{\tau}}\tau$ $3.27$
DM Observable Value DM Observable Value
$\Omega h^{2}$ $0.0091$ $\Omega h^{2}$ $0.02$
$\sigma^{\rm SI}_{p,n}/(10^{-48}\text{cm}^{2})$ $6.32,5.47$ $\sigma^{\rm SI}_{p,n}/(10^{-48}\text{cm}^{2})$ $1.78,4.85$
$\sigma^{\rm SD}_{p,n}/(10^{-43}\text{cm}^{2})$ $1.64,1.26$ $\sigma^{\rm SD}_{p,n}/(10^{-43}\text{cm}^{2})$ $4.08,3.13$
Table 3: Benchmark points P3 and P4 that satisfy all experimental constraints. Points P3 and P4 represent Higgsino-dominated DM in the $h_{1}$ and $h_{2}$ scenarios, respectively. All mass parameters are given in GeV.
3.3 Dark matter physics
To further illustrate the features of the GNMSSM parameter space in the light Higgs scenario, DM phenomenology was investigated. The package MicrOMEGAs  [81, 82, 83, 84, 85, 86, 87] was used to calculate the DM relic density and the spin-dependent (SD) and spin-independent (SI) DM-nucleon cross sections. It was assumed that a large DM population existed in the early universe and froze out to match the current Planck observation, $\Omega_{\rm DM}h^{2}=0.12\pm 0.01$ [88]. In this study, the relic density of DM (the lightest neutralino $\tilde{\chi}_{1}^{0}$) was required to be below the central value of 0.12. The SI and SD cross sections were scaled by a factor of $\Omega h^{2}/0.12$. It was found that in the $h_{1}$ scenario, the DM can be either Singlino- or Higgsino-dominated, whereas in the $h_{2}$ scenario, it is Higgsino-dominated.
Four benchmark points were selected, as listed in Tables 2 and 3, based on their dominant annihilation processes. These points also satisfy the 2024 LZ  [30] direct detection constraints. Points P1 and P2 correspond to Singlino-dominated DM in the $h_{1}$ scenario. After applying the LZ-2024 constraints, the values of $\lambda$ are generally small. For Singlino-dominated DM, $m_{\tilde{\chi}_{1}^{0}}\simeq m_{N}$, consistent with Eq.(2.17). For benchmark point P1, mostly annihilation channels of DM are $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to b\bar{b}/\tau^{+}\tau^{-}$. In contrast, for point P2, since $2m_{\tilde{\chi}_{1}^{0}}>m_{h_{s}}+m_{A_{s}}$, the channel $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}A_{s}$ is kinematically allowed and mostly annihilation channel of DM. This process occurs via s-channel exchange of CP-odd Higgs bosons and t-channel exchange of neutralinos. Additionally, the channel $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to A_{s}A_{s}$ is also open, proceeding through s-channel CP-even Higgs exchange and t-channel neutralino exchange.
Points P3 and P4 correspond to Higgsino-dominated DM in the $h_{1}$ and $h_{2}$ scenarios, respectively. For these benchmark points, mostly annihilation channels of DM are achieved through co-annihilation processes $\tilde{\chi}_{i}^{0}\tilde{\chi}_{1}^{\pm}\to XY$ ($i=1,2$), where X and Y represent quarks or leptons in SM, as well as through the annihilation channels $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to W^{+}W^{-}/ZZ$. In the Higgsino-dominated DM scenario, the masses of the lightest neutralino, the next-to-lightest neutralino, and the lightest chargino are nearly degenerate. Given the complexity of DM physics and its close relation to Higgs phenomenology, a detailed study of DM properties, particularly in the light Higgs scenario, will be the focus of future work.
@@ 4 Conclusion
This study systematically investigated the exotic decay channels of the standard-model-like Higgs boson into pairs of the lightest CP-odd or CP-even Higgs bosons within the framework of GNMSSM. First, we performed a comprehensive scan of the GNMSSM parameter space using the MultiNest algorithm, subject to experimental constraints from HiggsSignals-2.6.2 and HiggsBounds-5.10.2. To assess the implications of recent Higgs precision measurements for GNMSSM scenarios featuring light Higgs states (considering both cases where $h_{1}=h_{\text{SM}}$ and $h_{2}=h_{\text{SM}}$), we separately examined the exclusion limits provided by HiggsSignals and the HiggsBounds implementation in HiggsTools. Within the phenomenologically allowed parameter region, we further conducted a preliminary analysis of dark matter properties. The main findings were summarized as follows:
@@
  • In both the $h_{1}$ and $h_{2}$ scenarios, HiggsBounds imposes the most stringent constraints on the theoretical parameter space. In contrast, HiggsSignals and individual ATLAS searches exhibit comparatively weaker constraining power; moreover, the vast majority of parameter points excluded by these analyses are also disfavored by HiggsBounds . This is because HiggsBounds provides direct constraints from dedicated searches for non-Standard Model Higgs bosons, whereas HiggsSignals derives indirect constraints based on global compatibility with experimental data.
@@
  • It is worth noting that the analysis of the $h_{2}$ scenario reveals unique phenomenological features. Specifically, HiggsSignals can exclude certain parameter points with relatively low values for certain exotic decay branches, especially when $Br(h\to a_{1}a_{1}\to\tau\tau bb)\leq 2.5\%$. This exclusion effect stems from the kinematic accessibility of the $h_{2}\to h_{1}h_{1}$ decay process, which has a significantly enhanced branching ratio. Such enhanced decays will modify the properties of the standard model-like Higgs boson, thereby increasing the sensitivity of HiggsSignals to indirect constraints in regions where they were previously suppressed due to kinematic factors.
@@
  • Under the combined constraints of HiggsTools, both $h_{1}$ and $h_{2}$ scenarios require that the observed $h$ boson is predominantly SM-like ($V_{h}^{\rm SM}\geq 0.93$) with minimal singlet admixture ($V_{h}^{\rm S}\leq 0.32$). In the $h_{2}$ scenario, the lighter $h_{s}$ state must exhibit high singlet purity ($V_{h_{s}}^{\rm S}\geq 0.94$) and negligible SM-like components ($V_{h_{s}}^{\rm SM}\leq 0.32$).
  • In the $h_{1}$ scenario, DM can be either Singlino- or Higgsino-dominated, whereas in the $h_{2}$ scenario, it is Higgsino-dominated. For Higgsino-dominated DM, mostly annihilation channels of DM are achieved through co-annihilation with chargino; For Singlino-dominated DM, mostly annihilation channels of DM may be $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}A_{s}$.
Despite constraints from current Higgs data on light Higgs scenarios, phenomenologically viable regions in the parameter space persist. These regions merit further investigation for two principal reasons. First, dark matter physics exhibits considerable complexity and is deeply intertwined with Higgs phenomenology. Second, the corresponding mass spectrum can induce a strong first-order electroweak phase transition (SFOEWPT) in the early universe [89, 90]. This mechanism satisfies the necessary conditions for electroweak baryogenesis (EWBG), providing a plausible explanation for the observed baryon asymmetry of the universe, and simultaneously produces detectable gravitational wave (GW) signals during the phase transition. These aspects will form the core focus of our future work.